OpenStudy (anonymous):

what is the inverse of y=x^2+3

OpenStudy (anonymous):

is square root of -x+3 the same as the square root of x-3?

OpenStudy (across):

You have\[y=x^2+3.\] 1. Switch the variables.\[x=y^2+3\]2. Solve for \(y\).\[y=\pm\sqrt{x-3}.\]

OpenStudy (anonymous):

why is it positive & negative?

OpenStudy (anonymous):

When you take the square root of something, you have to take into consideration both the positive and negative.

OpenStudy (anonymous):

@across what about the inverse of y= -x^2+5

OpenStudy (anonymous):

The concept is the same process; switch it. So x = -y^2 + 5. so x-5 = -y^2. But simply divide by a negative sign so you get -x+5 = y^2. So y = sqrt(-x+5).

OpenStudy (anonymous):

*+/- squrt. Sorry about that.

OpenStudy (anonymous):

wait so its positive and negative too?

OpenStudy (anonymous):

@ilikedtea680 what about this one